# DeepUniUSTransformer: Towards A Universal UltraSound Model with Prompted Guidance

### v1

@@ 3 Experimental Results
Qualitative Analysis. We further employed t-SNE to visualize the distribution of representations. Specifically, we selected the breast datasets BUS-BRA, BUSIS, and UDIAT. The input to t-SNE was the features learned by the encoders of the single, fusion, and prompt models. As shown in Figure 3. There is clearly a noticeable domain shift in the single model result. In contrast, the fusion model results demonstrate a reduction in domain offset, with the datasets appearing more concentrated in the visualization. This indicates that the fusion model effectively mitigated the domain shift issue. Furthermore, the introduction of prompt messages resulted in nearly eliminating the domain offset. These findings align with the zero-shot results presented in Table 3. Finally, we visualized the segmentation results, as shown in Figure 4. The prompt model clearly outperformed the other models. This improvement can be attributed to the effective utilization of nature prompts and position prompts. By informing the model about the specific structure nature and its position, the prompt model achieved a deeper understanding of the segmentation task.
Figure 4: Some examples of segmentation result.Each column From left to right: original image, single, fusion, prompt, ground truth. The arrows show the details of how the edges are processed by different methods.
4 Conclusion
In this paper, we propose a universal model tailored for medical ultrasound imaging. Our model incorporates four distinct prompts, introducing information from various perspectives such as the input image nature, anatomical position, clinical task, and input image type. This comprehensive approach enhances the model’s learning capabilities. We conducted experiments using BroadUS-9.7K , a publicly available ultrasound dataset that we specifically collected. The results not only validate the effectiveness of our proposed framework but also demonstrate its potential for zero-shot testing. By expanding our analysis to include a wider range of datasets in the future, we aim to gain deeper insights into the effectiveness and adaptability of our proposed universal ultrasound model.

### v2

@@ III-B Comparison with SOTA and Ablation Study
III-C Evaluating Adapter Performance on New Datasets
TABLE III: Adapter performance comparison on BUSI dataset.
Dataset Task Single
UniUSNet
w/o prompt
UniUSNet Scratch Adapter
BUSI seg 60.85% 66.90% 66.91% 68.89% 69.56%
BUSI cls 80.80% 77.60% 78.40% 85.90% 86.40%
Average 70.83% 72.25% 72.66% 77.40% 77.98%
Our model allows for easy adjustment of the adapter layer to adapt to new datasets. Therefore, we conducted this experiment to explore the generalization ability of our model on new datasets. For this experiment, we selected the BUSI dataset from the BoardUS-9.7K dataset and excluded it from the training of the Single, UniUSNet w/o prompt, and UniUSNet experiments. We then performed zero-shot inference using the trained Single, UniUSNet w/o prompt, and UniUSNet models directly on the BUSI dataset and observed the results. (Note that since BUSI was not included in the Single experiment, we actually used the Single model obtained from the BUS-BRA dataset for zero-shot testing, as BUS-BRA is also a breast dataset and belongs to the same domain). To highlight the comparison, we also included an experimental setup named Scratch. In this setup, the BUSI dataset was used as the training set, and the results were obtained through training and testing. The Adapter setup involved freezing our model and fine-tuning the fully connected (FC) layer on the BUSI dataset.
As shown in Table III. The table revealed that the zero-shot performance of the Single model is unsatisfactory. The limited capacity and capability of a single model hinder its generalization performance. However, when comparing the UniUSNet method to UniUSNet w/o prompt, and UniUSNet w/o prompt to Single results, there is an observable improvement in model performance. This indicates that models trained on a large number of datasets possess a certain level of generalization ability, and the effectiveness of prompt incorporation. It is noteworthy that even with just 50 epochs of fine-tuned training based on the prompt architecture, the Adapter results surpass the Scratch performance achieved from training the model from scratch. This showcases the flexibility of our model and its ability to adapt to new datasets with minimal resources required for fine-tuning training.
IV Conclusion
In this paper, we propose a universal model tailored for medical ultrasound imaging. Our model incorporates four distinct prompts, introducing information from various perspectives i.e. the input image nature, anatomical position, clinical task, and input image type. This comprehensive approach enhances the model’s learning capabilities. We conducted experiments using BroadUS-9.7K , a publicly available ultrasound dataset that we specifically collected. The results not only validate the effectiveness of our proposed framework but also demonstrate its potential for adapter fine-tuning for new dataset. By expanding our analysis to include a wider range of datasets in the future, we aim to gain deeper insights into the effectiveness and adaptability of our proposed universal ultrasound model.
Acknowledgment
This work was supported by Science and Technology Development Fund of Macao (0021/2022/AGJ) and Science and Technology Development Fund of Macao (0041/2023/RIB2).